OpenStudy (anonymous):

Find the equation(s) of the tangent line(s) to the parabola y=x^2 throught the given point: (a) (0,a) and (b) (a,0)

OpenStudy (anonymous):

Well.. It's pretty hard to understand it the way you wrote it, but this is how I understand it: Assuming we have the following function: \[ f(x) = x^2 \] We have to points \[ a \to (0,a) \\ b \to (b, 0) \] to find a and b we have to use the other part of the coordinates \[ a \to (0,a) \\ f(0) = 0^0 = 0\\ a \to (0,0) \;\\\;\\ b \to(b,0)\\ f(x) = 0 \\ x^2 = 0 \\ x = 0 \\ b \to (0, 0) \] Since they both are same point then Using derivative \[ f'(x) = 2x \\ f'(0) = 2 \cdot 0 = 0 \] We have a slope 0 tangent line at (0,0) which makes it \[ y = 0 \cdot x + 0 = 0 \] Hope I got it right